Suppose you want to accumulate $10,000 for a down payment for a house. You will deposit $400 at the beginning of every month in an account that credits interest monthly at the rate of 0.6% per month. How long will it take you to achieve your goal?
Answer #1

As per given condition

400 * (F/A, 0.6%,n) = 10000

(F/A, 0.6%,n) = 10000 / 400 = 25

((1 + 0.006)^n-1)/0.006 = 25

((1.006)^n-1) = 25 * 0.006 = 0.15

1.006^n = 1.15

n = log 1.15 / log 1.006 = 23.36 months ~ 24 months